OpenStudy (anonymous):

Please walk me through this! 1. When bromine liquid is reacted with a solution of potassium iodide, a solution of potassium bromide and an aqueous solution of iodine result. a) Write a balanced chemical equation for this reaction. b) 30.5mL of liquid bromine (density=3.12g/mL) is reacted with a solution that contains 190.0g of potassium iodide. i) How many mols of bromine are reacted? Support with calculations. ii) How many mols of potassium iodide are reacted? Support with calculations. iii) Determine the limiting reactant. Support your choice with calculations.

OpenStudy (anonymous):

a) Br2 + 2 KI = 2 KBr + I2 b) (30.5 mL Br2) x (3.12 g/mL) / (159.808 g Br2/mol) = 0.595 mol Br2 (190.0 g KI) / (166.00277 g KI/mol) = 1.145 mol KI 1.14456 mole of KI would react completely with 1.145 x (1/2) = 0.573 mole of Br2, but there is more Br2 present than that, so Br2 is in excess and KI is the limiting reactant. So all of the KI would react.